# Coherent Audio-Visual Editing via Conditional Audio Generation Following Video Edits

### v1

@@ 1 Introduction
To address these challenges, we propose a new pipeline for joint audio-visual editing. As shown in Fig. 1, our method first utilizes advanced video editing techniques to create the target video, and then employs a novel video-to-audio generation model, which is conditioned on the source audio, target video, and a text prompt to edit the audio in alignment with the visual changes. We enhance the model architecture of the latest video-to-audio model [cheng2025mmaudio] to accept conditional audio input and introduce a data augmentation strategy for the new conditional input to improve training efficiency. Importantly, our model dynamically adjusts the influence of the source audio based on the complexity of the edits, maximizing preservation of the original audio structure. This approach leverages the strengths of state-of-the-art video editing while ensuring high audio-visual alignment. Experimental results demonstrate that our pipeline achieves superior performance compared to existing methods, particularly in maintaining audio-visual alignment and content integrity.
@@ 2.1 Flow matching
$\tilde{u}_{\theta}(\mathbf{x}_{t},t,\mathbf{C})=\$$u_{\theta}(\mathbf{x}_{t},t,\varnothing)$
$+w(u_{\theta}(\mathbf{x}_{t},t,\mathbf{C})-u_{\theta}(\mathbf{x}_{t},t,\varnothing)),$ (3)
where $\varnothing$ represents the null condition, and $w$ is a coefficient to control the strength of the guidance. To enable the model to handle the null condition, the conditional input is randomly dropped with a small probability (e.g., 0.1) during training.
@@ 2.2 Single-modal data editing
Audio editing methods can be broadly categorized into training-based and training-free approaches. In the training-based approach [wang2023audit, lan2025guiding], models are trained as conditional generators that take source audio as input and produce edited outputs. While conceptually straightforward, this method faces significant challenges in assembling large-scale datasets of paired source and target audio samples. Our method is also training-based one, but we avoid this issue by adopting acoustic feature extraction as described in Section 3.2.
In contrast, training-free approaches [manor2024zero, xu2024prompt, jia2025audioeditor] utilize pre-trained text-to-audio generation models [liu2023audioldm, liu2024audioldm, ghosal2023text] to generate the desired audio edits. To ensure preservation of unrelated content, these methods inject source audio information into the generation process using inversion techniques [song2021denoising] and attention control [hertz2023prompt]. Although this approach eliminates the need for training with specialized datasets, it typically incurs higher computational costs during generation.
@@
Training video generation models is considerably more resource-intensive than audio models. As a result, most video editing methods rely on pre-trained generative models. There are two main types of pre-trained models used: text-to-image and text-to-video models.
Early work leveraged text-to-image models such as Stable Diffusion [rombach2022high] due to their accessibility and ease of use. Since these models are only capable of generating single frames, specialized techniques [cong2024flatten, li2024vidtome, cohen2024slicedit, geyer2024tokenflow, kara2024rave] and fine-tuning strategies [wu2023tune, liu2024video] have been developed to improve temporal consistency across video frames.
Recently, the emergence of advanced text-to-video generation models [blattmann2023stable, yang2025cogvideox, wan2025wan] has enabled more effective video editing. Studies utilizing these models [fan2024videoshop, shi2024bivdiff, zhang2025v2edit, jiang2025vace] demonstrate excellent temporal consistency.
@@ 2.3 Joint audio-video editing
liang2024language introduced a lightweight module that embeds the source audio-video pair into a specialized text token. This token is then input into pre-trained text-to-image and text-to-audio models to generate the corresponding video frames and audio. In contrast, lin2025zero extended score distillation sampling to the cross-modal setting by jointly leveraging pre-trained text-to-image and text-to-audio models. They further incorporated a contrastive loss on relevant and irrelevant regions within the audio-video pair to improve consistency between audio and video edits.
@@
In our method, we address these limitations with a sequential approach to joint editing. First, we apply a state-of-the-art video editing technique to generate the target video, ensuring high temporal consistency. Next, we use our novel video-to-audio model, which incorporates the source audio as a conditional input, to conduct audio editing following the video edits, thereby achieving strong audio-visual alignment in much higher frame-rate videos (20 fps in the experiments).
@@ 3.1 Overview
Our goal is to build a model that generates edited audio aligned with the target edited video while preserving the acoustic structure of the source audio. To achieve this, we first extract hierarchical acoustic features from the source audio to capture its structural information at multiple levels of detail (Section 3.2). These features, together with conditional features from the target video and optionally a target prompt, are fed into the model (Section 3.3), which is trained to generate the edited audio using a flow matching framework. To improve training, we introduce a new data-augmentation strategy for the acoustic features called detail-temporal masking (Section 3.4). During generation, the model automatically selects the appropriate level of detail for the acoustic features based on the estimated editability score, which is determined by measuring the semantic similarity between the source audio and the target video (Section 3.5).
@@ 3.2 Hierarchical acoustic feature extraction
A common choice to extract audio features would be using a pretrained model. However, such a model generally does not satisfy the second property, as they are typically trained to be sensitive to differences in audio semantics (e.g., CLAP [elizalde2023clap]). In addition, extracting semantic features at a specified level of detail is challenging, which conflicts with the third property. Consequently, we design simpler statistical features of audio signals that satisfy all three properties. Statistical features, such as loudness, are less sensitive to semantic changes in audio content, making them more reliable for preserving structural information [garcia2025sketch2sound].
Figure 2: Hierarchical acoustic features.
@@
$\mathbf{a}_{l}=\mathrm{concat}\left(\{A(\mathbf{s}_{i}^{(l)})\}_{i=0,...,2^{l}-1}\right),$ (4)
@@
$\mathbf{a}=\mathrm{concat}\left(\{\mathbf{a}_{l}\odot\mathbf{m}_{l},\mathbf{m}_{l}\}_{l=0,...,L}\right),$ (5)
where $L$ is the maximum level, and $\mathbf{m}_{l}\in\mathbb{R}^{2^{l}\times T_{a}}$ denote a mask at level $l$, whose elements are all zero if $l$ is larger than the specified level, and all one otherwise. In the remainder of this paper, we refer to $\mathbf{a}$, consisting of the hierarchical loudness features and their corresponding mask (which specifies the level of detail), as the acoustic features.
@@ 3.3 Model architecture
Our model is based on the MMAudio architecture [cheng2025mmaudio], with two key modifications to effectively handle information from the acoustic features during the generation process.
@@
The backbone of the base model is a multi-modal Transformer that processes noisy audio latent features, as well as conditional text and video features, to predict the velocity for audio generation through the flow matching framework. Following the latest text-to-image models [flux2024], it comprises a sequence of MM-DiT blocks [esser2024scaling] followed by a sequence of audio-only single-modal DiT blocks. Additionally, the base model includes a conditional synchronization module to enhance audio-visual synchrony. This module extracts high frame-rate features from the input video using Synchformer [iashin2024synchformer] and combines them with temporally-global conditional features extracted from text and video. The combined features are fed into the DiT blocks through adaptive layer normalization [perez2018film].
@@
  • Modulation of audio latents: We add the processed acoustic features to the audio latents. Before the addition, we adjust the temporal length of the acoustic features by linear interpolation and process them with a trainable linear layer to match the dimensionality of the audio latents. This approach, which uses input addition or concatenation, is a common method for integrating conditional structural data into flow matching models and has been widely used in prior studies such as wang2024frieren and zhu2024flowie.
  • Modulation of Synchformer features: We also modulate the Synchformer features in a frame-wise manner using the acoustic features, which are processed with temporal length regulation and a feed-forward module (as used in the DiT block). This temporally-local modulation is important because the Synchformer features are used for adaptive layer normalization in the base model and thus strongly influence the generated audio in terms of temporal dynamics. To properly reflect the conditional acoustic information in the generation process, it is essential to modulate the Synchformer features before they are used in adaptive normalization.
@@ 3.4 Training with detail-temporal masking
  • Detail masking. For each training sample, we randomly determine the level of detail of the acoustic features. Since this corresponds to randomly choosing a proper mask for the acoustic features as described in Section 3.2, we refer to this augmentation as detail masking. The detail masking enables the model to work with various levels of details, which means that we can control how much the original acoustic structure in the source audio should be preserved in the target audio at the inference time. We also include a choice to totally mask out the entire acoustic features in the setting, allowing us to apply classifier-free guidance to our model for more accurate structure preservation.
  • Temporal masking. We also apply random masking along with the temporal dimension to prevent the model from excessively relying on the acoustic features to generate the target audio. Since the acoustic features contain temporally rich information of the audio, especially when the level of detail is high, the model tends to focus more on aligning with the acoustic features rather than text and video. This leads to degrading text fidelity and audio-visual alignment, which is not desirable for the audio editing task. To alleviate this issue, we mask out the acoustic features at randomly selected time frames during training. It enforces the model to jointly reflect both the acoustic features and the other conditional features to generate the target audio.
@@ 3.5 Generation with adaptive conditioning
Figure 5: The adaptive conditioning based on the editability score.
@@
$\tilde{u}_{\theta}(\mathbf{x}_{t},t,\mathbf{C},\mathbf{a})$ $=u_{\theta}(\mathbf{x}_{t},t,\varnothing,\varnothing)$
$\ +w_{1}(u_{\theta}(\mathbf{x}_{t},t,\mathbf{C},\varnothing)-u_{\theta}(\mathbf{x}_{t},t,\varnothing,\varnothing))$
@@
The optimal level of detail for the acoustic features depends on how difficult the audio editing task is. For easy edits, where the source audio is already similar to the target (e.g., bird chirping $\rightarrow$ seagull chirping), preserving detailed acoustic features helps maintain structural fidelity without sacrificing alignment with the target prompt and video. In contrast, for difficult edits, where significant changes are needed (e.g., cat meowing $\rightarrow$ lion roaring), retaining too much detail from the source is less effective, much of the original acoustic structure will be altered or lost. Therefore, our method adaptively adjusts the level of acoustic detail based on the estimated editing difficulty for each sample, aiming to balance structural preservation and target fidelity.
@@
where $s_{\mathrm{min}}$ and $s_{\mathrm{max}}$ are the expected maximum and minimum scores, respectively, and $l_{\mathrm{max}}$ denote the maximum level of detail. For example, an easy edit (high similarity, $s$ close to $s_{\mathrm{max}}$) results in a higher level of detail being preserved, while a difficult edit (low similarity, $s$ close to $s_{\mathrm{min}}$) results in less detail being retained.
We control the strength of adaptive conditioning by changing the value of $l_{\mathrm{max}}$. Allowing a larger level of detail leads to more accurate preservation of the original audio structure. On the other hand, we fix both $s_{\mathrm{max}}$ and $s_{\mathrm{min}}$. To determine these values, we empirically investigated the editability scores of genuine pairs of audio and video as well as those of random pairs in the VGGSound validation dataset [chen2020vggsound], which can be considered as the maximum and minimum levels of the editability score for this dataset. Their average scores are $0.32$ and $0.02$, and these are used to set $s_{\mathrm{max}}$ and $s_{\mathrm{min}}$, respectively.
@@ 4.1 Dataset and evaluation metrics
Following the prior work [cheng2025mmaudio], we jointly used several text-video-audio and text-audio datasets to train our model: VGGSound [chen2020vggsound], AudioCaps [kim2019audiocaps], Clotho [drossos2020clotho], and WavCaps [mei2024wavcaps]. All audio clips were resampled to 16 kHz, and 8-second segments were cropped without overlap for each training sample.
@@
  • Structure preservation (LPAPS) assesses how much the acoustic structure of the source audio is preserved in the target audio. Following the prior audio-editing studies [manor2024zero, lin2025zero], we used LPAPS [iashin2021taming] to evaluate the structure preservation.
  • Text fidelity (IB-TA) assesses the fidelity of the target audio to the given target prompt using ImageBind. Since ImageBind can only handle two-second audio clips, we cropped five non-overlapping audio sub-clips from the target audio and extracted ImageBind features from each sub-clip. Then, we computed the cosine similarity between those features and the feature extracted from the target prompt. To aggregate the multiple similarities, we used max pooling, because the audio event described in the target prompt is often temporally local (e.g., dog barking) and thus could be missed in several sub-clips.
@@ 4.2 Setup
We evaluated several methods representing two main approaches to audio-visual editing: independent and sequential.
@@
In this approach, audio and video are edited separately. For audio editing, we employed SDEdit [meng2022sdedit] and ZETA [manor2024zero], both adapted to use Stable Audio Open [evans2025stable], a state-of-the-art audio generation model. For video editing, we used RAVE [kara2024rave], which is based on Stable Diffusion 1.5 [rombach2022high], and VACE [jiang2025vace], which utilizes the latest video generation model, Wan 2.1 [wan2025wan]. Among the various editing tasks supported by VACE, we used the “structure transfer” in our experiments. To control the structure preservation in SDEdit and ZETA, we varied the intermediate timestep to inject the source audio information.
@@
In this approach, audio and video are edited in sequence. Our proposed method belongs to this category: we apply it to videos processed by RAVE and VACE and report the results (RAVE/VACE $\rightarrow$ Ours). We varied the value of $l_{\mathrm{max}}$ in Eq.7 to control the structure preservation. We also evaluated our method with fully masked acoustic features, which completely remove source audio information from conditioning. This setting corresponds to applying a naive video-to-audio model to the edited videos (RAVE/VACE $\rightarrow$ V2A).

### v2

@@ 1 Introduction
To address these challenges, we decouple joint audio-video editing into two sequential sub-tasks as shown in Fig. 1. We first apply an off-the-shelf video editing method to transform the source video into the target video, leveraging recent progress in high frame-rate video editing without modification. We then introduce a new task setting, audio editing following video edits, where the goal is to edit the source audio to align with the edited video while preserving the original audio structure as much as possible (e.g., the timing and number of sound events, and the continuity of background ambience), guided by a text prompt. This task is particularly challenging because the model must jointly satisfy text fidelity, audio-visual alignment to the edited visual events, and audio structure preservation.
For this new task, we propose a conditional video-to-audio editing model built by minimally extending a state-of-the-art video-to-audio generation model [cheng2025mmaudio]. Specifically, we modify the model to accept the source audio as an additional condition alongside the target video and text, enabling it to preserve the original audio structure. We further introduce a data augmentation strategy for this new conditional input to improve training efficiency, and we also adopt an adaptive conditioning mechanism to adjust the influence of the source audio based on the complexity of the edits, maximizing audio structure preservation. With these changes, our model can dynamically balance text fidelity, audio-visual alignment, and audio structure preservation. Experimental results demonstrate that our pipeline achieves superior performance compared to existing methods, particularly in maintaining audio-visual alignment and content integrity.
@@ 2.1 Flow matching
$\tilde{u}_{\theta}(\mathbf{x}_{t},t,\mathbf{C})=\$$u_{\theta}(\mathbf{x}_{t},t,\varnothing)+w(u_{\theta}(\mathbf{x}_{t},t,\mathbf{C})-u_{\theta}(\mathbf{x}_{t},t,\varnothing)),$ (3)
where $\varnothing$ denotes the null condition, and $w$ is a coefficient to control the strength of the guidance. To enable the model to handle the null condition, the conditional input is randomly dropped with a small probability (e.g., 0.1) during training.
@@ 2.2 Single-modal data editing
Audio editing methods can be broadly categorized into training-based and training-free approaches. In the training-based approach [wang2023audit, lan2025guiding], models are trained as conditional generators that take source audio as input and produce edited outputs. This approach is conceptually straightforward, but it requires large-scale datasets of paired source and target audio samples, which is challenging in practice. In contrast, training-free approaches [manor2024zero, xu2024prompt, jia2025audioeditor] leverage pre-trained text-to-audio generation models [liu2023audioldm, liu2024audioldm, ghosal2023text] to generate the desired audio edits. Although they eliminate the need for paired training datasets, they require techniques to inject source audio information into the generation process, which typically incurs significantly higher computational costs during generation. Our method is also training-based one, but we mitigate the dataset issue by adopting acoustic feature extraction as described in Section 3.2.
@@
Training video generation models from scratch is considerably more resource-intensive than audio models. As a result, most video editing methods rely on pre-trained generative models. There are two main types of pre-trained models used: text-to-image and text-to-video models.
Early work leveraged text-to-image models such as Stable Diffusion [rombach2022high] due to their accessibility and ease of use. Since these models are only capable of generating single frames, specialized techniques [cong2024flatten, li2024vidtome, cohen2024slicedit, geyer2024tokenflow, kara2024rave] and fine-tuning strategies [wu2023tune, liu2024video] have been developed to improve temporal consistency across video frames. Recently, the emergence of advanced text-to-video generation models [blattmann2023stable, yang2025cogvideox, wan2025wan] has enabled more effective video editing. Studies utilizing these models [fan2024videoshop, shi2024bivdiff, zhang2025v2edit, jiang2025vace] demonstrate excellent temporal consistency.
@@ 2.3 Joint audio-video editing
Liang et al. [liang2024language] introduced a lightweight module that embeds the source audio-video pair into a specialized text token. This token is then input into pre-trained text-to-image and text-to-audio models to generate the corresponding video frames and audio. In contrast, Lin et al. [lin2025zero] extended score distillation sampling to the cross-modal setting by jointly leveraging pre-trained text-to-image and text-to-audio models. They further incorporated a contrastive loss on relevant and irrelevant regions within the audio-video pair to improve consistency between audio and video edits.
@@
We address these limitations with a sequential approach to joint editing. First, we apply a state-of-the-art video editing technique to generate the target video, ensuring high temporal consistency. Next, we use our novel video-to-audio model, which incorporates the source audio as an additional conditional input, to conduct audio editing following the video edits, thereby achieving strong audio-visual alignment in much higher frame-rate videos (20 fps in the experiments).
@@ 3.1 Overview
Our goal is to build a model that generates edited audio aligned with the target edited video while preserving the acoustic structure of the source audio. To achieve this, we first extract hierarchical acoustic features from the source audio to capture its structural information at multiple levels of detail (Section 3.2). These features, together with conditional features from the target video and a target prompt, are fed into the model (Section 3.3), which is trained to generate the edited audio using a flow matching framework. To improve training, we introduce a new data-augmentation strategy for the acoustic features called detail-temporal masking (Section 3.4). During generation, the model automatically selects the appropriate level of detail for the acoustic features based on the estimated editability score, which is determined by measuring the semantic similarity between the source audio and the target video (Section 3.5).
@@ 3.2 Hierarchical acoustic feature extraction
A common choice to extract audio features would be using a pretrained model. However, such a model generally does not satisfy the second property, as they are typically trained to be sensitive to differences in audio semantics (e.g., CLAP [elizalde2023clap]). As a result, when used as “structure” features, they tend to overly preserve the original semantic content of the source audio (e.g., specific sound identities), which hinders semantic edits and can cause the editing to fail. In addition, extracting semantic features at a specified level of detail is challenging, which conflicts with the third property. Consequently, we design simpler statistical features of audio signals that satisfy all three properties. Statistical features, such as loudness, are less sensitive to semantic changes in audio content, making them more reliable for preserving structural information [garcia2025sketch2sound].
@@
$\mathbf{a}_{l}=\mathrm{concat}\left(\left\{A(\mathbf{s}_{i}^{(l)})\right\}_{i=0}^{2^{l}-1}\right),$ (4)
@@
$\mathbf{a}=\mathrm{concat}\left(\{\mathbf{a}_{l}\odot\mathbf{m}_{l},\mathbf{m}_{l}\}_{l=0}^{L}\right),$ (5)
where $L$ is the maximum level, and $\mathbf{m}_{l}\in\mathbb{R}^{2^{l}\times T_{a}}$ denote a mask at level $l$, whose elements are all zero if $l$ is larger than the specified level, and all one otherwise. In the remainder of this paper, we refer to $\mathbf{a}$ as the acoustic features.
@@ 3.3 Model architecture
Our model shown in Fig. 3 is based on the MMAudio architecture [cheng2025mmaudio], with two key modifications to effectively handle information from the acoustic features.
@@
The backbone of the base model is a multi-modal Transformer that processes noisy audio latent features, as well as conditional text and video features, to predict the velocity for audio generation through the flow matching framework. It comprises a sequence of MM-DiT blocks [esser2024scaling] followed by a sequence of audio-only single-modal DiT blocks. Additionally, the base model includes a conditional synchronization module to enhance audio-visual synchrony. This module extracts high frame-rate features from the input video using Synchformer [iashin2024synchformer] and combines them with temporally-global conditional features extracted from text and video. The combined features are fed into the DiT blocks through adaptive layer normalization [perez2018film].
@@
  • Modulation of audio latents: We add the processed acoustic features to the audio latents. Before the addition, we adjust the temporal length of the acoustic features by linear interpolation and process them with a trainable linear layer to match the dimensionality of the audio latents. This approach, which uses input addition or concatenation, is a common method for integrating conditional structural data into flow matching models and has been widely used in prior studies [wang2024frieren, zhu2024flowie].
  • Modulation of Synchformer features: We also modulate the Synchformer features in a frame-wise manner using the acoustic features, which are processed with temporal length regulation and a feed-forward module (as in the DiT block). This temporally-local modulation is important because the Synchformer features strongly influence the generated audio in terms of temporal dynamics via adaptive layer normalization. To properly reflect the conditional acoustic information in the generation process, it is essential to modulate the Synchformer features before they are used in the base model.
@@ 3.4 Training with detail-temporal masking
Figure 5: The adaptive conditioning based on the editability score.
@@
  • Detail masking. For each training sample, we randomly determine the level of detail of the acoustic features. Since this corresponds to randomly choosing a proper mask for the acoustic features as described in Section 3.2, we refer to this augmentation as detail masking. This masking enables the model to work with various levels of details, which means that we can control how much the original acoustic structure in the source audio should be preserved in the target audio at the inference time. We also include a choice to totally mask out the entire acoustic features in the setting, enabling classifier-free guidance for more accurate structure preservation.
  • Temporal masking. We also apply random masking along with the temporal dimension to prevent the model from excessively relying on the acoustic features to generate the target audio. Since the acoustic features contain temporally rich information of the audio, especially when the level of detail is high, the model often focus more on aligning with the acoustic features rather than text and video. This leads to degrading text fidelity and audio-visual alignment, which is not desirable for the audio editing. To alleviate this issue, we mask out the acoustic features at randomly selected time frames during training. It enforces the model to jointly reflect both the acoustic features and the other conditional features to generate the target audio.
@@ 3.5 Generation with adaptive conditioning
$\tilde{u}_{\theta}(\mathbf{x}_{t},t,\mathbf{C},\mathbf{a})=u_{\theta}(\mathbf{x}_{t},t,\varnothing,\varnothing)$ $+w_{1}(u_{\theta}(\mathbf{x}_{t},t,\mathbf{C},\varnothing)-u_{\theta}(\mathbf{x}_{t},t,\varnothing,\varnothing))$
@@
The optimal level of detail depends on how difficult the audio editing task is. For easy edits, where the source audio is already similar to the target (e.g., bird chirping $\rightarrow$ seagull chirping), preserving detailed acoustic features helps maintain structural fidelity without sacrificing alignment with the target prompt and video. In contrast, for difficult edits, where significant changes are needed (e.g., cat meowing $\rightarrow$ lion roaring), retaining too much detail from the source is less effective, much of the original acoustic structure will be altered or lost. Therefore, our method adaptively adjusts the level of acoustic detail based on the estimated editing difficulty for each sample, aiming to balance structural preservation and target fidelity.
@@
where $s_{\mathrm{min}}$ and $s_{\mathrm{max}}$ are the expected maximum and minimum scores, respectively, and $l_{\mathrm{max}}$ denote the maximum level of detail. For example, an easy edit ($s$ close to $s_{\mathrm{max}}$) results in a higher level of detail being preserved, while a difficult edit ($s$ close to $s_{\mathrm{min}}$) results in less detail being retained.
We control the structure preservation in our model by changing the value of $l_{\mathrm{max}}$. Allowing a larger level of detail leads to more accurate preservation of the original audio structure. On the other hand, we fix both $s_{\mathrm{max}}$ and $s_{\mathrm{min}}$. To determine these values, we empirically investigated the editability scores of genuine pairs of audio and video as well as those of random pairs in the VGGSound validation dataset [chen2020vggsound], which can be considered as the maximum and minimum levels of the editability score for this dataset. Their average scores are $0.32$ and $0.02$, and these are used to set $s_{\mathrm{max}}$ and $s_{\mathrm{min}}$, respectively.
@@ 4.1 Dataset and evaluation metrics
Following MMAudio [cheng2025mmaudio], we jointly used several text-video-audio and text-audio datasets to train our model: VGGSound [chen2020vggsound], AudioCaps [kim2019audiocaps], Clotho [drossos2020clotho], and WavCaps [mei2024wavcaps]. All audio clips were resampled to 16 kHz, and 8-second segments were cropped without overlap for each training sample.
@@
  • Structure preservation (LPAPS) assesses how much the acoustic structure of the source audio is preserved in the target audio. Following the prior studies [manor2024zero, lin2025zero], we used LPAPS [iashin2021taming] to evaluate the structure preservation.
  • Text fidelity (IB-TA) assesses the fidelity of the target audio to the given target prompt using ImageBind. Since ImageBind can only handle two-second audio clips, we cropped five non-overlapping audio sub-clips from the target audio and extracted ImageBind features from each sub-clip. Then, we computed the cosine similarity between those features and the feature extracted from the prompt. To aggregate the multiple similarities, we used max pooling, because the audio event described in the prompt is often temporally local (e.g., dog barking) and thus could be missed in several sub-clips.
@@ 4.2 Setup
We evaluated several methods representing three main approaches to audio-visual editing: independent, sequential, and joint.
@@
In this approach, audio and video are edited separately. For audio editing, we employed SDEdit [meng2022sdedit] and ZETA [manor2024zero], both adapted to work with Stable Audio Open [evans2025stable], a state-of-the-art audio generation model. For video editing, we used RAVE [kara2024rave], which is based on Stable Diffusion 1.5 [rombach2022high], and VACE [jiang2025vace], which utilizes the latest video generation model, Wan 2.1 [wan2025wan]. Among the various editing tasks supported by VACE, we used the “structure transfer" setting. To control the structure preservation in SDEdit and ZETA, we varied the intermediate timestep at which we inject the source audio information.
@@
In this approach, audio and video are edited in sequence. Our proposed method belongs to this category: we apply it to videos processed by RAVE and VACE (RAVE/VACE $\rightarrow$ Ours). We also evaluated our method with fully masked acoustic features, which completely remove source audio information from conditioning. This setting corresponds to applying a naive video-to-audio model to the edited videos (RAVE/VACE $\rightarrow$ V2A).
Joint approach.
In this approach, audio and video are edited jointly. We chose AvED [lin2025zero] as the state-of-the-art baseline. We controlled the structure preservation by varying the guidance scale used in the score distillation.